OpenStudy (anonymous):

soybean meat is 18% protein cornmeal is 9% protein .How many pound of each should be mixed together in order to get 360-lb mixture that is 16% protein

OpenStudy (anonymous):

Suppose s is the number of pounds of soy, then c=360-s would be the number of pounds of corn, where c represents pounds of corn. Then, multiply the concentration times the amount for each commodity, and add them up and set equal to the total: \[.18s+.09(360-s)=.16(360)\] If you solve this for s you can then subtract that number from 360 to get the pounds of corn.

OpenStudy (anonymous):

Solving you get \[.18s+32.4-0.9s=57.6\] \[0.9s=25.2\] \[s=280\] To get the pounds of corn \[c=360-280=80\]

OpenStudy (anonymous):

so this would be 80 for cornmeal